What nominal rate of​ interest, compounded​ monthly, corresponds to an effective rate of 11​%?
Answer by addingup(3677)

0.11 = (1+(i/n))^(n)-1
i = interest rate
n = number of compounding periods
0.11 = (1+(i/12)^12-1 = 10.48 nominal rate.